Can someone help me solve with dimensional analysis and steps like how I’m doing it please. I uploaded 39-41 but I am stuck on the others
md of total parenteral nutrition (TPN) is ordered to be infused from 1900-0700. At what r should the IV pump be programed? 29. 3000 m2 pf total min should 40. Mr. Jones is to be infsas pepic lrdisease, has Pepcid 20 mg in 100 mL DsW ordered q12 h The a to be infused over 30 be given? minutes With a tubing drop factor of 10 gtt/mL, how many gtts/min s Amalnourished patient has an order for 500 ml of Intralipid 10% to be infused over 6 hours. At what rate should the IV pump be programmed? 41. mL hy 42. An order is received to infuse penicillin G 4,000,000 units in 100 mL of DsW q12 h. The tubing factor is 10 gtt/mL. The penicillin should be infused over 45 minutes. How many gtts/min should be given? 43. Mrs. Sanchez has hypokalemia. The physician orders potassium 60 mEq in 250 ml of DsW. The facilitys policy states to infuse the potassium at 20 mEq/h. At what rate should the IV pump be programmed? 44. A postop patient has Kefzol 1 g ordered q8 h. The Kefzol is dissolved in s0 mL of DsW and is to be infused over 35 minutes. The tubing factor is 60 gtt/mL. How many gtts/min should be given? 45. After a total hip replacement, a patient has an order for Toradol 60 mg IVPB q6 h over 15 minutes The Toradol is diluted in 25 mL of DsW. The tubing drop factor is 15 gtt/mL. How many gtts/min should be given?
hy hw hain min
Answer #1

ANSWER

42. Calculation of flow are (gtts/min) of pencillin given.

                        Formula

                            Flow rate (gtts/min) = Volume(ml) / Time(min) *drop factor.(gtt/ml)

                               Volume =100 ml.

                                 Time = 45 min.

                            Drop factor = 10 gtt/ml

                       Flow rate = 100 /45*10

                                           = 2.2 * 10

                                           = 22.2

                                           = 22 gtts/min.

43. Flow rate of pottasium 60 mEq in a IV pump.

                      Formula

                             Flow rate = Volume(ml) / time(hr)

                               Volume = 250 ml.

                                 Time = 20 mEg / hr

                                          = 60 mEq / 3 hr.

                                           = 3 hr.

                Flow rate = 250 / 3

                                 = 83.3

                                 = 83 ml /hr.

44.Flow rate (gtts/min) of pottasium to be given in 35 minutes.

                Formula

                     Flow rate (gtt/min) = Volume(ml) / time(min) * drop factor(gtt/ml)

                              Volume = 50 ml.

                               Time = 35 min.

                               Drop factor = 60 gtt/ml

                    Flow rate = 50 / 35 * 60

                                       = 1.43 *60

                                       = 85.8

                                      = 86 gtts/min.

45. Flow rate (gtts/min) of Toradol to be given in 15 minutes.

                    Formula

                             Flow rate = Volume(ml) / time(min) * drop factor(gtt/ml).

                             Volume = 25 ml.

                              Time = 15 mt.

                              Drop factor = 15 gtt / ml.

                           Flow rate = 25 / 15 * 15

                                            = 1.67 * 15

                                           = 25 gtts/ min.